2. Suppose that X1, X2, ..., Xn N(41,01) and Yı,Y2,...,Ym * N(H2;02) are two independent random samples. (a) What is E[X -
Answer #1

Sop: Xi , Xxj. .· ,Xn HoN(A,, G;?). and Y,%,.., Ym A NI Ab, G?) • ) X1 , Xx,. , Xn and Y, Y, , . »,Ym are andependent IEEIXVar (x-Ÿ) = Now, 675 = stder.(x-9) = Var lx-Ÿ) ² O77 = Stderl 8-7) = =-> © Greven that , 6,²=2; 6²= 2.5 6= 2,5 n=10, m = 15 x> P[ 1*-7 - (44-42) Z 0.5] PL IZI so.826] $ (0.826) - I fo.826) ewhere $) Cs the cdif. f zNN = O (0.826) – ( 1- 0 c0.621) O-@Geiren that; 0,² = 2 , 6,2 = 2,5 and we need to find Amallest value af n such that PE IX-7-(4-14)|so5],0.95 S0.5 2+2.5 4.5 Oand O•95 -1- K2 = 0.05 K2 Too (00 = 20 0•25. K² = 0.5 -> 20 = 0.25n Tos 4. 45*2* J = 4.5* P0 0:25 n n = 45 * 8 = 360 o: m-n=3